# The Geometry behind the Two-Dimensional Hard-Disk ‘Glass Transition’

### v1

@@ Abstract
The relation between dynamics and structure in systems of bidisperse 2D hard disks with apparent arrested dynamics is considered using numerical simulations and Voronoi analysis. Surprisingly, the suspensions systematically appear to fall out of equilibrium at an area fraction of $\phi\approx 0.777$ over a wide range of disk-size ratios. This is in close agreement with the experimental findings of [Lozano et al., Nat. Mater. 18, 1118 (2019)] for a single large-to-small size ratio of $\approx 1.4$. Even inside the crystalline region of the state diagram there are weak structural signatures present for this area fraction. Adopting a granocentric viewpoint, this allows for the identification of a geometric feature — a floret pentagonal tiling — that could underlie the observations. That is, this tiling has an area fraction of $\phi_{g}=\sqrt{3}\pi/7\approx 0.777343$, which closely matches the apparent ‘glass transition’ value of $\phi$. The connection suggests that a zero-entropy, ground-state tiling can influence the dynamics of the system at finite temperature, in a manner that may be similar to the way in which geometry in foam-like tissue models induces a jamming transition. A geometric ground-state argument also appears to explain the change in local dynamics observed by [Li et al., Nature 587, 225 (2020)], when a honeycomb lattice is used as the reference. Combined these findings change how to view high-$\phi$ 2D fluids, namely as being influenced by the presence of zero-entropy geometric ground states. The proposed connection leads to an unexpected (dynamic) transition in the finite-temperature system, which can be mistaken for a(n onset) glass transition, but is in a class of its own. There is also a tentative connection to random loose packing for frictional systems. An outlook is provided on how this concept may be extended to soft-particle systems and into higher dimensions.
I Introduction
@@
Arrested dynamics in amorphous soft-matter systems has been a topic of significant interest for decades [3, 4, 5, 6, 7, 8, 9]. The experimental realization of colloidal glasses made up of (almost) hard spheres [10, 11, 12], has given the field a strong impetus to consider the more fundamental question on the nature of the ‘glass transition’  [^1] Here, I am purposefully circumspect in using the word ‘transition’, since the formation of a glass is unlike well-known transitions such as freezing or melting. The latter two involve discontinuities in derivatives of the free energy and are therefore readily identified as true transitions. The formation of a glass, however, appears to follow a smooth increase in the viscosity of dozens (if not more) orders of magnitude. This also means that the system typically falls out of equilibrium and the free energy cannot be properly defined.. In hard-particle systems, the glass transition takes place at a density for which the system exhibits a relaxation-time divergence  [^2] The more commonly referenced quantity of temperature, i.e., glass-transition temperature, is not particularly relevant for hard-sphere systems, as it merely scales the unit of time., without its structure changing significantly from that of a fluid [15]. Despite the kinetic nature of this transition and a lack of obvious concurrent structural change, there has been a dedicated effort into obtaining structural signatures in experiment [16, 17, 18, 19], simulation [20, 21, 22, 23, 24, 25], and theory [26, 27, 28].
@@ I.1 Experimental Inspiration and Open Question
Recently, however, an experimental study [1] revealed that self-propelled probe particles give a strong signal close to the ‘glass transition’ of hard spheres confined to a quasi-2D environment. That is, the orientation of the self-propelled probes is such that their rotational dynamics may be tracked and this changes upon varying the area fraction $\phi$ of the bidisperse hard-sphere mixture that surrounds it. The probe’s rotational diffusion $D_{p}$ was found to be strongly enhanced with respect to its free ($\phi=0$) value close to where the translational diffusion in the passive system drops off $\phi=\phi_{g}\approx 0.776$; even for small self-propulsion speeds. In addition, $D_{p}$ was found to be peaked at $\phi_{g}$ and to decay asymmetrically away from its peak value as a function of $\phi$. In Ref. [1], this observation was related to the increase in the bulk relaxation time close to their ‘glass transition,’ using a theoretical model that incorporated a time-correlated reorientational noise based on the relaxation time.
@@
It is important to note that in the experiments of Lozano et al. [1], the $\alpha$ relaxation times obtained from the self-intermediate scattering function follow the Vogel-Fulcher-Tammann (VTF) law [58, 27] exceptionally well. The VFT fit in Ref. [1] indicates a divergence at $\phi\approx 0.845$, which is close to where random close packing (rcp) for this colloid size ratio ($R\approx 1/1.4$) sets in, see also the ESI [59]. This seems to suggest that there is no glass transition at $\phi_{g}$. Yet, the MSD curves in the experimental system do indicate a strong dynamical change close to $\phi_{g}$. This was related to the system falling out of equilibrium and, hence, where the ‘glass transition’ is located.
@@
Numerical modeling of the probe dynamics in the experimental system [60], performed in my group showed that an asymmetric peak $D_{p}$ can be obtained using a contact-based friction that induces reorientation. However, it proved difficult to place the peak value close to the experimentally observed $\phi_{g}\approx 0.776$. The $D_{p}$ peaks were systematically located closer to the $\phi$ value associated with rcp. We speculated that this may be attributed to a lack of dissipative interactions, such as contact friction or lubrication, that we did not account for in our modeling of the passive fluid. Thus, while our simple model was able to produce qualitatively the features observed in experiment, there remained an open question as to the nature of the ‘glass transition’ in this system.
@@
This study was followed up by an experimental and numerical work on a quasi-2D system of colloidal rods [61], which was initiated in the Bechinger group and where my group provided the numerical modeling. Together, we revealed that there is a $D_{p}$ enhancement for spherical self-propelled probes close to the glass transition, which is well captured using the reorientation model of Ref. [60]. In addition, we found that only considering the passive (thermal) rod particles is sufficient to accurately predict the enhancement, meaning that the probes pick up on local features present in the passive system alone. More intriguingly, there was a visually evident structural change in the system close to $\phi_{g}$, namely the breaking up of large rafts of colloidal rods into smaller clusters [61].
I.2 Dual Goal of the Study
Both the difficulty in reproducing numerically the experimental ‘glass transition’ in Ref. [60] and the presence of an apparent structural signal for enhancement in the rod-like systems [61], suggest that it is worth revisiting the quasi-2D hard-sphere system of Ref. [1]. One aim of this work is therefore to identify structural features that correlate with (the onset of) arrested dynamics and thereby understand the origin of the observed $\phi_{g}$.
@@ I.2 Dual Goal of the Study
Given the prevalence of Voronoi-based analysis in (effective) 2D modeling, I will follow a tessellation-based analysis route. However, I additionally draw inspiration from the adjacent field of epithelial tissue modeling, in particular, foam-like vertex- and Voronoi-based approaches [62, 63, 64, 65, 66, 67, 68, 69]. When there are $N$ cells, these systems typically possess a Hamiltonian of the form
@@
where $A_{i}$ and $P_{i}$ are the instantaneous area and perimeter length of the $i$-th (Voronoi) cell, respectively. The parameter $A_{0}$ is the target area and $P_{0}$ represents the target perimeter length, which are the rest ‘sizes’ for the harmonic terms in the Hamiltonian. Note that these target shape parameters are common to all cells.
@@
It is well-established that there is a transition from an amorphous fluid-like state to an amorphous solid-like state when the shape index $p_{0}\equiv P_{0}/\sqrt{A_{0}}=p_{0}^{\ast}\approx 3.81$ [62, 63]. The transition shape index $p_{0}^{\ast}$ is identical to that of a regular pentagon — a 5-sided polygon for which all edges and angles are equal. The physical intuition is that this choice of $p_{0}$ leads to frustration, since regular pentagons do not tesselate space. That is, there is a mismatch between the local requirement on the cells to be pentagonal and the global requirement on the model epithelium to be confluent (cover all of space).
Bi et al. [62] showed that the onset of arrested dynamics — here referred to as a jamming transition as the system need not be thermal  [^3] Technically, this only holds for vertex models in the zero-temperature limit [115]. — can be correlated to the appearance of a barrier that hinders topological exchange between neighboring cells. The second aim of this work is to extent these ideas to hard-particle systems and what insight they can provide into the structural origin of the reported ‘glass transition’ [1].
I.3 Approach and Result
@@
I circumvent the issues in previously modeling the experimental system using soft potentials [60] by studying numerically 2D systems of hard disks using event-driven molecular dynamics (EDMD) simulations [71]. I focus on a stoichiometric ratio of 50:50 small to large disks, as in the experiment [1], but I also analyze a far larger range of size ratios $R\in[1/1.7,1]$ and area fractions $\phi$. Here, $R$ is the ratio of the small to the large disk diameter. For these systems, I determine the radical Voronoi diagram [50] and show how to use the statistics of the particle neighborhoods to obtain a state diagram. I base these statistics on the isoperimetric quotient $q$, which is a variant of the aforementioned shape index $p$ ($q=4\pi/p^{2}$).
Using the mean of $q$, I can determine when the system crystallizes, which occurs when the average shape is that of a $5.5$-gon. I will introduce how to define the isoperimetric quotient of a polygon with a non-integer number of edges and show how this reveals a re-entrant amorphous state as $\phi$ is increased at a sufficiently low value of $R$. In fact, my state diagram appears to have a ‘critical point’ in $R$ and $\phi$, which correlates well with the system falling out of equilibrium. Although the analogy should not be taken too far, given that above the ‘glass transition’ the system is slow to equilibrate.
Surprisingly, I find that for $R\lesssim 0.83$ the area fraction, for which the particle dynamics appears to become arrested, is nearly constant $\phi\approx 0.777$. I confirm this measurement, which is based on the mean squared displacement (MSD), by examining features in a shell-based percolation measure. A small feature in the amount of pentagonal neighborhoods for the large particles in the systems tracks this arrested-dynamics line very well. Unexpectedly, this even extends the line $\phi\approx 0.777$ into the region where the system crystallizes, which includes monodisperse systems. I analyze the established literature on structure and dynamics in 2D systems in detail and find hints in several theory and simulation papers that support my finding of a constant-$\phi$ behavior.
The fact that non-equilibrium features are present in the otherwise crystalline monodisperse system, inspired me to adopt a granocentric viewpoint. That is, I consider the properties of a neighborhood around a single disk in an otherwise monodisperse system. This allows me to identify a convex pentagonal shape for the Voronoi cell around this disk, when it is in contact with 5 neighbors. The configuration has a local area fraction $\phi_{g}=\sqrt{3}\pi/7\approx 0.777343$. Interestingly, this shape may be assembled into a floret pentagonal tiling [72, 73], which is a known way to tile space with congruent pentagonal elements. In this particle organization, each disk is in direct contact with 5 neighbors, which makes it a local entropic minimum with both zero configurational and vibrational entropy.
I.4 Conclusion
The combined body of evidence that I present, allows me to propose an unexpected origin of the ‘glassy dynamics’ in the considered system. Namely, the pentagonal-floret tiling is a zero-entropy ‘geometric ground state’, wherein every particle is completely immobilized. The presence of this ground state echoes through in the disordered fluid, in a similar fashion as setting the shape index in a foam-like tissue model. That is, the floret pentagonal tiling enforces the slow dynamics at $\phi_{a}\approx 0.777$, thereby localizing what is identified as a ‘glass transition’ in the original experiment [1]. I also provide tentative evidence that the onset of caging in the system can be viewed in a similar fashion. I propose that it is related to a honeycomb-lattice geometric ground state and make the connection to the experimental work by Li et al. [2].
The crux to this understanding lies in moving away from the idea that (2D) fluids are disordered systems. Instead, at least at sufficiently high density, these should be considered as entangled with ordered low-entropy ground states that can enforce geometric frustration even at finite temperature. The rapid change in dynamics observed in experiment, as well as the response behavior when subjected to active and/or external probing [1, 2], is then not because of subtle changes in the structure of the fluid. Instead, it is caused by radical changes in the entropy of the entangling (reference) geometric ground state. I will also provide arguments to support this viewpoint, although within the scope of this work, I was unable to come to a full theoretical understanding.
I.5 Structure of the Paper

### v2

# The Geometry behind the Glass Transition and Frictional Jamming in Systems of Two-Dimensional Hard Disks

@@ Abstract
The relation between dynamics and structure in systems of Brownian bidisperse two-dimensional (2D) hard disks with arrested dynamics is examined using numerical simulations. Surprisingly, the suspensions show dynamic arrest at an area fraction of $\phi\approx 0.777$ over a wide range of disk-size ratios. This is in close agreement with the experimental findings of [Lozano et al., Nat. Mater. 18, 1118 (2019)] ($\phi\approx 0.776$) for a quasi-2D colloidal suspension of spheres with large-to-small size ratio of approximately $1.4$. Intriguingly, this also matches a jamming transition ($\phi\approx 0.773$ to $0.777$) found experimentally in a 2D bidisperse granular packing of disks for a similar aspect ratio [Naseer et al. & Vishali et al., Powders and Grains (2025)]. Adopting a geometric viewpoint allows for the identification of the floret pentagonal tiling ($\phi=\sqrt{3}\pi/7\approx 0.777343$), which is comprised of congruent (elongated) pentagonal tiles. In analogy to foam and tissue models, it is the presence of this congruent reference state that induces a dynamical transition in the disordered fluid of the disks. That is, the change in dynamics observed both in a uniformly compressed granular medium and a colloidal one at finite temperature are implied to be caused by the same topological mechanism. Extending this reasoning to a honeycomb lattice could also explain the experimentally observed onset of caging dynamics ($\phi\approx 0.6$) in a similar bidisperse, quasi-2D system of colloidal spheres [Li et al., Nature 587, 225 (2020)]. An outlook is provided on how this concept may be applied to three dimensions (3D): removing one in four particles from a face-centered-cubic arrangement leads to a square bipyramidal honeycomb with associated volume fraction $\eta=\pi/(4\sqrt{2})\approx 0.55536$. The ideas put forward here thus also provide a possible origin for random loose packing of (frictional) hard spheres in 3D and can potentially shed light onto the nature of the glass transition.
I Detailed Background
@@
Arrested dynamics in amorphous soft-matter systems has been a topic of significant interest for decades [5, 6, 7, 8, 9, 10, 11]. The experimental realization of colloidal glasses made up of (almost) hard spheres [12, 13, 14], has given the field a strong impetus to consider the more fundamental question on the nature of the glass transition  [^1] The formation of a glass is unlike well-known transitions such as freezing or melting. The latter two involve discontinuities in derivatives of the free energy and are therefore readily identified as true transitions. The formation of a glass, however, appears to follow a smooth increase in the viscosity of dozens (if not more) orders of magnitude. This also means that the system typically falls out of equilibrium and the free energy cannot be properly defined.. In hard-particle systems, the glass transition takes place at a density for which the system exhibits a relaxation-time divergence  [^2] The more commonly referenced quantity of temperature, i.e., glass-transition temperature, is not particularly relevant for hard-sphere systems, as it merely scales the unit of time., without its structure changing significantly from that of a fluid [17]. Despite the kinetic nature of this transition and a lack of obvious concurrent structural change, there has been a dedicated effort into obtaining structural signatures in experiment [18, 19, 20, 21, 22, 23], simulation [24, 25, 26, 27, 28, 29], and theory [30, 31, 32].
@@ I.1 Experimental Inspiration and Open Question
Recently, however, an experimental study [1] revealed that self-propelled probe particles give a strong signal close to the glass transition of hard spheres confined to a quasi-2D environment. That is, the orientation of the self-propelled probes was tracked based on the half-coating of their surface, from which the angular mean squared displacement (aMSD) was subsequently obtained. This was found to change upon varying the area fraction $\phi$ of the bidisperse hard-sphere mixture ($\approx 4.4/6.3\approx 0.698; RR$ is the ratio of the small-to-large disk size) that surrounds it. The probe’s aMSD proved linear for the time-differences considered, which allowed for the extraction of the rotational diffusion coefficient $D_{p}$, where the labelling stands for ‘probe’. The rotational diffusion coefficient proved to be strongly enhanced with respect to its free ($\phi=0$) value, close to where the translational diffusion in the passive system drops off at $\phi\equiv\phi_{g}\approx 0.776$; even for small self-propulsion speeds, see Fig. 1. In addition, $D_{p}$ decayed asymmetrically away from its peak value as a function of $\phi$. In Ref. [1], this observation was related to the increase in the bulk relaxation time close to their glass transition  [^3] The glass-transition area fraction was obtained experimentally by studying the behavior of the translational diffusion coefficient of the passive system. using a theoretical model that incorporated a time-correlated reorientational noise.
Figure 1: Experimental results for the active probe’s rotational diffusion. The effective rotational diffusion coefficient $D_{p}$ normalized by the bare rotational diffusion $D_{0,r}$ of a self-propelled probe as a function of the area fraction $\phi$ of the surrounding bidisperse passive spheres (close to a surface, hence quasi-2D). The black dots show the mean value and the black bars the standard deviation, see Ref. [1]. The two red curves are a stretched-exponential (left) and an exponential (right) fit to the data and serve to guide the eyes. The vertical blue dashed line indicates the value $\phi=\sqrt{3}\pi/7\approx 0.777$, which will be justified later. Data is reproduced with permission from the authors.
@@
It is important to note that in the experiments of Lozano et al. [1], the $\alpha$ relaxation times obtained from the self-intermediate scattering function follow the Vogel-Fulcher-Tammann (VTF) law [63, 31] exceptionally well. The VFT fit in Ref. [1] indicates a divergence at $\phi~\approx~0.845$, which is close to where random close packing (RCP) for this colloid size ratio sets in, see also the ESI [64]. This calls into question what the nature of the transition at $\phi_{g}\approx 0.776$ is. However, the mean squared displacement (MSD) curves in the experimental system indicate a strong reduction of the passive diffusion coefficient close to $\phi_{g}$. This was related to the system falling out of equilibrium and hence $\phi_{g}$ is where the glass transition is located in Ref. [1]. Here, I will therefore use the term ‘glass’ and ‘glass transition’ thoughout.
@@
Numerical modeling of the probe dynamics in the experimental system [65], performed in my group, showed that an asymmetric peak $D_{p}$ can be obtained using a contact-based friction that induces reorientation. This approach also accounted for the short-time nature of the ballistic component to the aMSD, which occured at times that were not experimentally resolved [1]. However, it proved difficult to place the peak value close to the experimentally observed $\phi_{g}\approx 0.776$. The simulated $D_{p}$ peaks were systematically located closer to the $\phi$ value associated with RCP. In Ref. [65], it was speculated that this may be attributed to a lack of dissipative interactions, such as contact friction or hydrodynamic lubrication, that was not accounted for in our modeling of the passive fluid. In view of the results presented below, the likely orgin is, however, an insufficiently hard-core interaction. A detailed explanation of the experimentally observed transition at $\phi_{g}$ was not provided.
@@
This study was followed up by an experimental and numerical work on a quasi-2D system of colloidal rods [66], which was initiated in the Bechinger group and where my group provided the numerical modeling. Together, it was revealed that there is a $D_{p}$ enhancement for spherical self-propelled probes close to the glass transition of the passive rod fluid, which is well captured using the reorientation model of Ref. [65]. Due to the slower dynamics of the rods, we were also able to probe the ballistic enhancement of the aMSD. In addition, considering only the passive (thermal) rod particles proved sufficient to accurately predict the enhancement, meaning that the probes pick up on local features present in the passive system alone. More intriguingly, there was a visually evident structural change in the system close to $\phi_{g}$, namely the breaking up of large rafts of colloidal rods into smaller clusters [66]. This suggested that a structural change could have been overlooked in the modeling [65] of the experimental system of quasi-2D fluids of hard spheres [1].
I.2 Granular Realization and Topological Insight
For athermal particles, i.e., ones for which Brownian motion is not expected to play a significant role, the jamming transition is considered instead of the glass transition [67, 68, 69]. This is a transition between flowing and rigid states of amorphous systems. Properties of the networks of contact forces that form in granular materials have been connected with the mechanical response of such systems, including the appearance of the jamming transition [70, 71, 72, 73]. It is further known that the viscosity $\eta$ of a granular suspension can be fitted using the Krieger-Dougherty relation [74] and that the nature of the forces between particles plays an important role in setting the area (or volume) fraction for which the shear-thickening curve diverges. Experiments [75], theory [76], and simulations [77, 78, 79, 80, 81] have shown a branched structure to the response, where the branches are related to the nature of the frictional contacts between particles. These branches control, for example, where RCP (frictionless) and random loose packing (RLP; fully frictional) appear in a hard-particle suspension.
Recent experimental works in 2D [82, 3, 2] reveal that the jamming transition in a frictional 2D bidisperse disk suspension ($\phi\approx 0.773$ - $0.777$) lies close to the where Lozano et al. [1] report a glass transition in a colloidal system. In particular, the authors of Ref. [3] performed a topological analysis on force networks obtained using bidisperse ($R\approx 11/15.4\approx 0.714$) photoelastic disks subjected to isotropic compression. Their measurements show a topological change around $\phi\approx 0.77$, with the number of homology generators in the persistence diagram of dimension-0 (connected components) saturating at $\phi\approx 0.7773$. This reflects that the isolated force networks merge, which the authors report as predictive of mechanical stability and an early marker of the jamming transition, as also found in earlier work on the topic [82]. Note that the granular experiments are athermal, while those performed in the Bechinger group [1, 66] are subject to Brownian dynamics. However, the former are performed on an air table and may therefore may have configurations closer to those found in a thermal system than other means of preparing jammed granular packings.
Generally, the jamming and glass transition are considered to have commonalities [83, 84], but the underlying connection, if present, is not yet fully established. The striking similarity in $\phi$ where the glass transition [1] and jamming transition [3] are found for two distinct 2D systems begs further examination. The topological nature of the jamming transition obtained in the latter, suggests that there could be a feature shared by both. This is because topological properties enjoy a measure of protection, even in finite-temperature systems.
I.3 Further Transitions in 2D Thermal Experiments
[…]
The displacement response of the system has similarities to the peak in the rotational diffusion of the self-propelled probe by Lozano et al. [1]. However, given the difference in area fraction and system response, their origin may be different. That is, the MSD data in Ref. [4] does not reveal a clear drop, as is expected for a glass transition. If such a transition is present in the Li et al. system, it may be that it occurs above $\phi>0.73$. Lozano et al. [1] did not reveal a convincing probe signal at the values of $\phi$ for which Li et al. find their displacement peak. It may be that the probe’s rotational diffusion coefficient starts to rise at around $\phi\approx 0.6$, but it is not possible to verify this based in the data in Ref. [1]. The frictional-reorientation model introduced in Ref. [65] shows some enhancement for all values of $\phi<\phi_{g}$, but it may not be appropriate for low area fractions.
I.4 Topological Transitions and Geometry in 2D
@@
The connection between topology / geometry and response in 2D transitions is elegantly shown in epithelial tissue modeling, and, in particular, in foam-like vertex- and Voronoi-based approaches [85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95]. When there are $N$ cells, these systems typically possess a Hamiltonian that contains terms of the form
@@
where $A_{i}$ and $P_{i}$ are the instantaneous area and perimeter length of the $i$-th (Voronoi) cell, respectively. The parameter $A_{0}$ is the target area and $P_{0}$ represents the target perimeter length, which are the rest ‘sizes’ for the harmonic terms in the Hamiltonian. Note that these are common to all cells; heterogeneity in these parameters has only recently been studied [96].
@@
It is well-established that, in the limit of vanishing cell activity, there is a transition from an amorphous fluid-like state to an amorphous solid-like state when the shape index $p_{0}\equiv P_{0}/\sqrt{A_{0}}$ assumes the critical value $p_{0}^{\ast}\approx 3.81$ [85, 86]. The transition value is identical to the $p_{0}$ belonging to a regular pentagon — a 5-sided polygon for which all edges and angles are equal. The physical interpretation is that this choice of $p_{0}$ leads to frustration, since regular pentagons do not tesselate space, while the foam-like models impose confluence. That is, there is a mismatch between the local requirement on the cells to be pentagonal and the global requirement on the model epithelium to cover all of space.
Figure 2: Visualization of topological transitions in a 2D system of disks. (a) Neighbor exchange of particles (red disks) is a T1 transition, as shown using part of the Voronoi lattice for this configuration (black lines). In the central plot, the particles are equidistantly placed and the transition takes place. (b) In a T2 transition, a particle is removed or added. This is illustrated here by shrinking one of the particles. In the middle row, the disk is smaller for every other particle and made increasingly opaque from left to right, starting from the second particle on the left. The black lines indicate the radical Voronoi diagram for this configuration.
Bi et al. [85] showed that the onset of arrested dynamics — here referred to as a jamming transition as the system need not be thermal  [^4] Technically, this only holds for vertex models in the zero-temperature limit [151]. — can be correlated to the appearance of a barrier that hinders topological exchange between neighboring cells. There is also a more obvious transition at the critical value of $p_{0}^{\dagger}=\sqrt{8\sqrt{3}}\approx 3.72$, which is where the system has hexagonal cells [98], and it seems likely that this can principle be extended to other space-filling tessellations. Importantly, the topological transition in this system is neighbor exchange, referred to as a T1 transition. There are also T2 transitions, which are associated with the introduction and removal of particles [99, 95], see Fig. 2 for a visual representation.
I.5 Goal of the Paper
Taking these separate pieces of evidence into consideration, here, I hypothesize that there is a topological feature that underlies both the glass transition in thermal and the frictional jamming transition (at RLP) in granular hard-disk systems in 2D. In addition, I further hypothesize that this topological transition has an elegant geometric interpretation, analogous to the one found in foam and tissue models. Potentially, this hypothesis can be extended into three dimensions (3D), for which I will provide preliminary results in this work.
@@ I.5 Goal of the Paper
To investigate the proposed connection, I will circumvent the issues in previously modeling the experimental system using soft potentials [65] by studying numerically 2D systems of hard disks using event-driven molecular dynamics (EDMD) simulations [100]. Here, I will focus on a stoichiometric ratio of 1:1 small to large disks, as in the experiment by Lozano et al. [1], but I will analyze a far larger range of size ratios $R\in[1/1.7,1]$ and area fractions $\phi$. This will allow me to determine whether there are lines of near-constant area fraction, for which transitions occur. The presence of these lines means that the particle aspect ratio is not relevant for the transition.
Combined with the experimental evidence that the transition is (potentially) independent of particle size, it becomes feasible to instead examine features of a single component. This simplifies the adoption of a granocentric viewpoint [101, 102, 103]. That is, one where the properties of a neighborhood around a single disk is considered, since this may then be done for a monodisperse system. I will propose a geometric interpretation via this route, which points toward a floret pentagonal tiling [104, 105]. That is, this tiling is a geometric ground state, which underlies the observed dynamic arrest in both thermal and athermal systems. This idea will subsequently be tested further, in order to show the relevance of geometric ground states to the response of 2D fluids of hard particles. The analysis reveals that the result is robust and that a honeycomb-lattice ground state (equilateral-triangle Voronoi tiling) can also impact the dynamics, potentially explaining the observation by Li et al. [4].
I.6 Structure of the Paper